OpenStudy (anonymous):

find the y-intercept for 4x-4y=-4

myininaya (myininaya):

set x=0 and solve for y

myininaya (myininaya):

4(0)-4y=-4 -4y=-4 y=1

OpenStudy (anonymous):

Yeah or you could rewrite it as y=x+1 by dividing both sides by 4 then adding y to bothe sides then adding 1 to both sides the answer is still 1 like myininaya said

OpenStudy (anonymous):

-4y=4x-4 y=-x+1 y int b=1